Consider the below circuit and suppose ω = 500 rad/seQuestion 15 1 pts Consider the below circuit and suppose w = 500rad/ sec. What is the equivalent ADMITTANCE with respect to tc. What is the equivalent ADMITTANCE with respect to terminals a and b?
Answer #1

Solution - del = 500 rad/ Sec = 4 mH c = 3MF Ye = = -0.54. Skemons 500 x4 x16 3 Y = juc Ye = j 500x 3 xlos ye = j 1500x1016 =You + Ye 73 rela As Yab = Yc+ YL Yab = j15 x 10-4+(-0.53) Yab = (j15 x 10^-.5j) Siement] Joy option D is right.